OpenStudy (anonymous):

A proton moves at a speed of 8x10^6 m/s  along the x-axis. Iteneters a region where there is a magnetic field of strength  2.5 T, directed at angle of 60 degrees to the x-axis and lying in the xy plane.  What is the force on the proton and acceleration?

OpenStudy (irishboy123):

you should start here: \(\mathbf{F} = q\left[\mathbf{E} + (\mathbf{v} \times \mathbf{B})\right]\), Lorentz Force Law  here you should find that: \(\mathbf{v} \times \mathbf{B} = |\mathbf{v}| \; |\mathbf{B}| \sin \frac{\pi}{3}\)